# SloPalSpeech: A 2,800-Hour Slovak Speech Corpus from Parliamentary Data

### v1

@@ 6.  Benchmarks
To evaluate the impact of parliamentary speech fine-tuning on Slovak ASR performance, we conducted a comprehensive benchmarking of all produced models. While preliminary indicators were available from the training process, final assessments were carried out on the official test sets of Common Voice 21 and FLEURS. As with all benchmarking we included basic text normalization before producing the WER value.
6.1.  Improvements
Model CV21 FLEURS $\Delta$ (%)
Small-Base 58.4 36.1 32.7-32.7 / 25.5-25.5
Small-Tuned 25.7 10.6
Medium-Base 38.0 18.7 20.0-20.0 / 11.1-11.1
Medium-Tuned 18.0 7.6
@@ 6.1.  Improvements
Table 5: WER comparison of Whisper Small and Medium (Base vs fine-tuned). $\Delta$ indicates absolute WER reduction (%) for CV21 / FLEURS.
Model CV21 FLEURS $\Delta$ (%)
Turbo-v3 Base 31.7 10.7 18.5-18.5 / 4.3-4.3
Turbo-v3 Tuned 13.2 6.4
Large-v3 Base 20.8 9.2 9.2-9.2 / 3.7-3.7
Large-v3 Tuned 11.6 5.5
Table 6: WER comparison of Whisper Large-v3 variants (Base vs fine-tuned). $\Delta$ indicates absolute WER reduction (%) for CV21 / FLEURS.
@@
The benchmarking outcomes, presented in Table 5 for Whisper Small and Medium and in Table 6 for Whisper Large-v3-Turbo and Large-v3, show WER decreases across all models, with the largest relative improvements observed in the smaller ones. For Whisper Small, fine-tuning reduced WER by approximately 65–70% compared to the baseline, approaching the base performance of Large-v3. This demonstrates that the fine-tuned dataset effectively adapted model weights, with diminishing marginal gains as model size increased. Manual inspection confirmed that fine-tuning improved both word-level accuracy and the generation of Slovak-like token compositions, even when predictions were partially incorrect. Whisper Medium showed similar but smaller improvements, still yielding substantial performance gains.
@@
Despite Whisper Large-v3 being trained on millions of hours of data, our dataset further reduced WER by a few percentage points. At these low error rates, improvements were less apparent during manual inspection, though handling of rare Slovak words noticeably improved. Among all evaluated models, Whisper Large-v3-Turbo offered the best performance-to-size trade-off: with roughly 730 million fewer parameters than Large-v3 and only about 100 million more than Medium, it achieved accuracy within about 1% of Large-v3, making it a strong candidate for efficient inference.
7.  Results
We compiled a large-scale dataset of 2.8k hours of Slovak parliamentary hearings paired with transcripts, suitable for ASR model training. Due to the long-form nature of the recordings and transcripts, our alignment method could not cover the entire collection. Therefore, in addition to the ASR dataset, we release all speaker-segmented transcripts, comprising approximately 60 million words.
To validate the dataset and develop accurate Slovak ASR models for real-world applications, we fine-tuned four Whisper models of varying sizes. Evaluated on Common Voice 21 and FLEURS, all models achieved reduced WER, with particularly large relative improvements for smaller models and notable gains for larger ones.
@@ 8.  Conclusion
We presented a complete pipeline for creating a large-scale ASR dataset, from sourcing and collecting raw data to alignment, segmentation, and demonstrating its effectiveness in training real-world models. With 2.8k hours of speech, this work helps move Slovak ASR beyond low-resource status. We further show the strong utility of parliamentary speech for Slovak ASR training and provide a methodology for aligning long-form audio with extensive transcripts.
We publish all resources, including the ASR dataset (2.8k hours [^3] This dataset includes only the portion of audio successfully aligned with transcripts. The full set of segmented transcripts (60 million words) is a superset of the text in the ASR dataset.), the complete collection of segmented transcripts (60 million words), and the fine-tuned Whisper models (small, medium, large-v3-turbo, and large-v3).
@@
Whisper models are known to hallucinate due to their architecture. While the base model often repeated words like "Ďakujem," our fine-tuned models occasionally produced sequences such as "Pán poslanec," reflecting a bias toward parliamentary speech. We recommend adopting an approach similar to Faster Whisper SYSTRAN (2025), where compression ratio checks during inference can trigger a re-decode with adjusted temperature to mitigate hallucinations. Additionally, when encountering unknown words or phrases, the model may substitute them with parliamentary terms. For example, in one Common Voice sample with the transcript "ostrov Nihau" the model predicted "ostrov Mihál" where "Mihál" is a former politician’s surname. We assume that similar cases are likely to happen.
Since the fine-tuning process updated all model parameters, the model shifted strongly toward Slovak. While we did not quantify the effect, we observed degraded English transcription performance, suggesting a loss of multilingual capability.
9.  Future Work
Parliamentary proceedings are transcribed in many European countries, where several languages remain low-resource for ASR. With access to recordings and full transcripts, our alignment method could be scaled to other languages to boost ASR resources beyond Slovak.
10.  Acknowledgements
@@
We thank VÚB Banka for their generous support and for providing the GPU resources necessary to accomplish this work.
11.  Bibliographical References
[…]
@@ 11.  Bibliographical References
  • Apache Software Foundation (2025) Apache Software Foundation. 2025. Apache tika project wiki. https://cwiki.apache.org/confluence/display/TIKA/Home.
@@
  • Ardila et al. (2019) Rosana Ardila, Megan Branson, Kelly Davis, Michael Henretty, Michael Kohler, Josh Meyer, Reuben Morais, Lindsay Saunders, Francis M Tyers, and Gregor Weber. 2019. Common voice: A massively-multilingual speech corpus. arXiv preprint arXiv:1912.06670.
@@
  • Bain et al. (2023) Max Bain, Jaesung Huh, Tengda Han, and Andrew Zisserman. 2023. Whisperx: Time-accurate speech transcription of long-form audio. arXiv preprint arXiv:2303.00747.
@@
  • Conneau et al. (2023) Alexis Conneau, Min Ma, Simran Khanuja, Yu Zhang, Vera Axelrod, Siddharth Dalmia, Jason Riesa, Clara Rivera, and Ankur Bapna. 2023. Fleurs: Few-shot learning evaluation of universal representations of speech. In 2022 IEEE Spoken Language Technology Workshop (SLT), pages 798–805. IEEE.
@@
  • Devlin et al. (2018) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018. BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. CoRR, abs/1810.04805. ArXiv:1810.04805.
  • FFmpeg Developers (2025) FFmpeg Developers. 2025. Ffmpeg documentation.
@@
  • Gandhi (2022) Sanchit Gandhi. 2022. Fine‑Tune Whisper For Multilingual ASR with Transformers.
@@
  • Gugger et al. (2022) Sylvain Gugger, Lysandre Debut, Thomas Wolf, Philipp Schmid, Zachary Mueller, Sourab Mangrulkar, Marc Sun, and Benjamin Bossan. 2022. Accelerate: Training and inference at scale made simple, efficient and adaptable. https://github.com/huggingface/accelerate.
@@
  • Huggingface (2024) Huggingface. 2024. Whisper large-v3 finetuning - forum post.
@@
  • Levenshtein (1966) Vladimir I. Levenshtein. 1966. Binary codes capable of correcting deletions, insertions and reversals. Soviet Physics Doklady, 10(8):707–710. Originally published in Russian in Doklady Akademii Nauk SSSR (1965).
@@
  • Lodagala (2023) Vasista Sai Lodagala. 2023. Fine-tuning and evaluating whisper models for automatic speech recognition.
@@
  • National Council of the Slovak Republic and Parliament of the Czech Republic (2002) National Council of the Slovak Republic and Parliament of the Czech Republic. 2002. Joint czech and slovak digital parliamentary library. https://www.nrsr.sk/dl/.
@@
  • (13) NRSR. Abecedný zoznam poslancov národnej rady slovenskej republiky. https://www.nrsr.sk/web/?sid=poslanci/zoznam_abc.
@@
  • (14) NRSR. Mediaportál národnej rady slovenskej republiky. https://tv.nrsr.sk/.
@@
  • Radford et al. (2023) Alec Radford, Jong Wook Kim, Tao Xu, Greg Brockman, Christine McLeavey, and Ilya Sutskever. 2023. Robust speech recognition via large-scale weak supervision. In International conference on machine learning, pages 28492–28518. PMLR.
@@
  • Srivastav et al. (2023) Vaibhav Srivastav, Somshubra Majumdar, Nithin Koluguri, Adel Moumen, Sanchit Gandhi, et al. 2023. Open automatic speech recognition leaderboard.
@@
  • SYSTRAN (2025) SYSTRAN. 2025. Systran/faster-whisper. GitHub repository.
@@
  • Wang et al. (2021) Changhan Wang, Morgane Riviere, Ann Lee, Anne Wu, Chaitanya Talnikar, Daniel Haziza, Mary Williamson, Juan Pino, and Emmanuel Dupoux. 2021. Voxpopuli: A large-scale multilingual speech corpus for representation learning, semi-supervised learning and interpretation. arXiv preprint arXiv:2101.00390.
@@
  • Wolf et al. (2020) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Rémi Louf, Morgan Funtowicz, Joe Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien Plu, Canwen Xu, Teven Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest, and Alexander M. Rush. 2020. Transformers: State-of-the-art natural language processing. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 38–45, Online. Association for Computational Linguistics.
@@
  • Zhao et al. (2023) Yanli Zhao, Andrew Gu, Rohan Varma, Liang Luo, Chien-Chin Huang, Min Xu, Less Wright, Hamid Shojanazeri, Myle Ott, Sam Shleifer, et al. 2023. Pytorch fsdp: experiences on scaling fully sharded data parallel. arXiv preprint arXiv:2304.11277.
  • Zusag et al. (2024) Mario Zusag, Laurin Wagner, and Bernhad Thallinger. 2024. Crisperwhisper: Accurate timestamps on verbatim speech transcriptions. In Proc. Interspeech 2024, pages 1265–1269.

### v2

@@ 6.  Benchmarks
To evaluate the impact of parliamentary speech fine-tuning on Slovak ASR performance, we conducted a comprehensive benchmarking of all produced models. While preliminary indicators were available from the training process, final assessments were carried out on the official test sets of Common Voice 21 and FLEURS. As with all benchmarking, we included basic text normalization before producing the WER value.
6.1.  Results
Model CV21 (%) FLEURS (%)
Small-Base 58.4 36.1
Small-Tuned 25.7 (32.7-32.7) 10.6 (25.5-25.5)
Medium-Base 38.0 18.7
Medium-Tuned 18.0 (20.0-20.0) 7.6 (11.1-11.1)
Turbo-v3 Base 31.7 10.7
Turbo-v3 Tuned 13.2 (18.5-18.5) 6.4 (4.3-4.3)
Large-v3 Base 20.8 9.2
Large-v3 Tuned 11.6 (9.2-9.2) 5.5 (3.7-3.7)
@@ 6.1.  Results
Table 5: WER comparison of Whisper models (Base vs fine-tuned). Improvements are shown in parentheses as absolute WER reduction in percentage points.
@@
The benchmarking outcomes on the whole test sets of Common Voice 21 and FLEURS, presented in Table 5, show WER decreases across all models after fine-tuning, with the largest relative improvements observed for the smaller models. Fine-tuning Whisper Small reduced WER by 65–70%, bringing a 244M-parameter model within striking distance of the 1.5B-parameter base Large-v3. Gains diminish with model size, indicating that approximately 2,800 hours of in-domain data most benefits smaller, deployment-friendly models. Manual inspection confirmed that fine-tuning improved both word-level accuracy and the generation of Slovak-like token compositions, even when predictions were partially incorrect. Whisper Medium showed similar but smaller improvements, still yielding substantial performance gains.
@@
Despite Whisper Large-v3 being trained on millions of hours of data, our dataset further reduced WER by a few percentage points. At these low error rates, improvements were less apparent during manual inspection, though handling of rare Slovak words noticeably improved. Among all evaluated models, Whisper Large-v3-Turbo offers the best efficiency trade-off: it matches Large-v3 accuracy within 1 percentage point while using 730M fewer parameters, making it the recommended choice for production Slovak ASR deployment.
Table 6 compares our largest fine-tuned model with NVIDIA’s Canary-1B-v2 and Parakeet-TDT-0.6B-v3 Sekoyan et al. (2025). We observe improvements over Canary on both benchmarks. Note that FLEURS and Common Voice v7 were included in the NVIDIA models’ training data, giving them an advantage on these test sets.
Model CV21 (%) FLEURS (%)
canary-1b-v2 14.46 6.82
Large-v3 Tuned 11.6 (2.86-2.86) 5.5 (1.32-1.32)
parakeet-v3 8.83 10.07
Large-v3 Tuned 11.6 (+2.77) 5.5 (4.57-4.57)
Table 6: WER comparison of our fine-tuned Whisper Large-v3 versus NVIDIA Canary and Parakeet models on Slovak benchmarks. Differences shown in parentheses relative to the baseline model (green = our model better, red = baseline better). Note: FLEURS and Common Voice v7 were included in NVIDIA models’ training data.
7.  Released Resources
We present SloPal, a comprehensive Slovak parliamentary resource with three integrated components:
SloPal Text Corpus: 330,000 speaker-segmented transcripts (66M words, 220M tokens) spanning 2001–2024, providing coverage of all successfully retrieved parliamentary proceedings. Each segment includes speaker metadata (first name, last name, parliamentary role), session information (number, date, type), and transcriber notes. This structure enables applications beyond ASR: political discourse analysis, speaker-specific language modeling, diachronic linguistic studies, and parliamentary rhetoric research. The comprehensive metadata and speaker-level segmentation distinguish SloPal as the most comprehensive Slovak parliamentary text resource.
SloPalSpeech Audio Dataset: 2,806 hours of aligned audio-transcript pairs with 30-second segments optimized for Whisper training. While audio coverage (68.5%) is lower than EuroSpeech’s 89%, the complete textual record remains available in SloPal.
[…]
Licensing and availability.
The source parliamentary recordings and transcripts are public government documents with no copyright restrictions. We release the SloPal text corpus and SloPalSpeech aligned audio under the CC BY 4.0 license, and the fine-tuned Whisper models under the same license as the original Whisper models (MIT). All resources are available on Hugging Face. [^5] https://huggingface.co/collections/NaiveNeuron/slopal
@@ 8.  Conclusion
[…]
We release all resources publicly: SloPal text corpus (330,000 segments, 66M words), SloPalSpeech aligned audio (2,806 hours), and four fine-tuned Whisper models [^6] The text corpus includes all successfully parsed transcripts. SloPalSpeech contains the subset (2,806 hours) where audio was successfully aligned., providing comprehensive resources for Slovak NLP, ASR, and computational social science research.
9.  Future Work
Parliamentary proceedings in dozens of national legislatures worldwide offer transcribed speech in low-resource languages. We plan to apply our anchor-based alignment pipeline to additional Slavic languages where parliamentary corpora exist but lack Whisper-optimized alignment. We also intend to investigate parameter-efficient fine-tuning (e.g., LoRA) to preserve multilingual capability while adapting to individual languages. We encourage other research groups to replicate our pipeline for their target languages using the publicly available code and data.
@@ 9.  Future Work
Whisper models are known to hallucinate due to their architecture. While the base model often repeated words like "Ďakujem," our fine-tuned models occasionally produced sequences such as "Pán poslanec," reflecting a bias toward parliamentary speech. We recommend adopting an approach similar to Faster Whisper SYSTRAN (2025), where compression ratio checks during inference can trigger a re-decode with adjusted temperature to mitigate hallucinations. Additionally, when encountering unknown words or phrases, the model may substitute them with parliamentary terms. For example, in one Common Voice sample with the transcript "ostrov Nihau" the model predicted "ostrov Mihál" where "Mihál" is a former politician’s surname. Similar substitutions are expected in domains with vocabulary outside the parliamentary register.
Since fine-tuning updated all model parameters, the models specialize toward Slovak at the cost of other languages—a known trade-off of full-parameter fine-tuning. Parameter-efficient methods such as LoRA could preserve multilingual capability, which we leave to future work.
Acknowledgements
@@
We thank VÚB Banka for their generous support and for providing the GPU resources necessary to accomplish this work. This research was partially funded by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I02-03-V01-00029.
Bibliographical References
@@
  • Apache Software Foundation (2025) Note: https://cwiki.apache.org/confluence/display/TIKA/Home External Links: Link Cited by: §3.4.
@@
  • R. Ardila, M. Branson, K. Davis, M. Henretty, M. Kohler, J. Meyer, R. Morais, L. Saunders, F. M. Tyers, and G. Weber (2019) Common voice: a massively-multilingual speech corpus. arXiv preprint arXiv:1912.06670. Cited by: §3.
@@
  • M. Bain, J. Huh, T. Han, and A. Zisserman (2023) Whisperx: time-accurate speech transcription of long-form audio. arXiv preprint arXiv:2303.00747. Cited by: §4.1, footnote 2.
@@
  • A. Conneau, M. Ma, S. Khanuja, Y. Zhang, V. Axelrod, S. Dalmia, J. Riesa, C. Rivera, and A. Bapna (2023) Fleurs: few-shot learning evaluation of universal representations of speech. In 2022 IEEE Spoken Language Technology Workshop (SLT), pp. 798–805. Cited by: §3.
@@
  • J. Devlin, M. Chang, K. Lee, and K. Toutanova (2018) BERT: Pre-training of Deep Bidirectional Transformers for Language Understanding. CoRR abs/1810.04805. Note: arXiv:1810.04805 Cited by: §5.1.
  • T. Erjavec, M. Ogrodniczuk, P. Osenova, N. Ljubesic, K. Simov, A. Pancur, M. Rudolf, M. Kopp, S. Barkarson, S. Steingrímsson, Ç. Çöltekin, J. de Does, K. Depuydt, T. Agnoloni, G. Venturi, M. C. Pérez, L. D. de Macedo, C. Navarretta, G. Luxardo, M. Coole, P. Rayson, V. Morkevicius, T. Krilavicius, R. Dargis, O. Ring, R. van Heusden, M. Marx, and D. Fiser (2023) The ParlaMint corpora of parliamentary proceedings. Language Resources and Evaluation 57 (1), pp. 415–448. External Links: Document Cited by: §1.1.
  • FFmpeg Developers (2025) External Links: Link Cited by: §3.3.
@@
  • S. Gandhi (2022) Fine‑Tune Whisper For Multilingual ASR with Transformers. External Links: Link Cited by: §5.2.
@@
  • S. Gugger, L. Debut, T. Wolf, P. Schmid, Z. Mueller, S. Mangrulkar, M. Sun, and B. Bossan (2022) Accelerate: training and inference at scale made simple, efficient and adaptable.. Note: https://github.com/huggingface/accelerate Cited by: §5.3.1.
  • I. R. Helgadóttir, R. Kjaran, A. B. Nikulásdóttir, and J. Guðnason (2017) Building an ASR corpus using Althingi’s parliamentary speeches. In Proceedings of Interspeech 2017, pp. 2163–2167. External Links: Document Cited by: §1.1, Table 2.
@@
  • Huggingface (2024) Whisper large-v3 finetuning - forum post. External Links: Link Cited by: §5.3.2.
  • B. Kulebi, C. Armentano-Oller, C. Rodriguez-Penagos, and M. Villegas (2022) ParlamentParla: a speech corpus of Catalan parliamentary sessions. In Proceedings of the Workshop ParlaCLARIN III within the 13th Language Resources and Evaluation Conference, Marseille, France, pp. 125–130. Cited by: §1.1, Table 2.
@@
  • V. I. Levenshtein (1966) Binary codes capable of correcting deletions, insertions and reversals. Soviet Physics Doklady 10 (8), pp. 707–710. Note: Originally published in Russian in Doklady Akademii Nauk SSSR (1965) Cited by: §4.2.
  • Y. Liu, X. Yang, and D. Qu (2024) Exploration of Whisper fine-tuning strategies for low-resource ASR. EURASIP Journal on Audio, Speech, and Music Processing 2024 (1), pp. 14. Cited by: §1.
  • N. Ljubešić, P. Rupnik, and D. Korˇzinek (2024) The ParlaSpeech collection of automatically generated speech and text datasets from parliamentary proceedings. In arXiv preprint arXiv:2409.15397, Cited by: §1.1, Table 2, Table 2, Table 2.
@@
  • V. S. Lodagala (2023) Fine-tuning and evaluating whisper models for automatic speech recognition. External Links: Link Cited by: §5.2.
@@
  • National Council of the Slovak Republic and Parliament of the Czech Republic (2002) Note: https://www.nrsr.sk/dl/ External Links: Link Cited by: §3.1.
@@
  • [18] NRSR Abecedný zoznam poslancov národnej rady slovenskej republiky. Note: https://www.nrsr.sk/web/?sid=poslanci/zoznam_abc Cited by: §3.4.
@@
  • [19] NRSR MediaPortál národnej rady slovenskej republiky. Note: https://tv.nrsr.sk/ Cited by: §3.1.
  • S. Pfisterer, F. Grötschla, L. A. Lanzendörfer, F. Yan, and R. Wattenhofer (2025) EuroSpeech: a multilingual speech corpus. External Links: 2510.00514, Link Cited by: §1.2, Table 2.
@@
  • A. Radford, J. W. Kim, T. Xu, G. Brockman, C. McLeavey, and I. Sutskever (2023) Robust speech recognition via large-scale weak supervision. In International conference on machine learning, pp. 28492–28518. Cited by: §1, §1, §5.1.
  • F. Rekathati (2023) RixVox: a Swedish speech corpus with 5500 hours of speech from parliamentary debates. Note: Published by KBLab, National Library of Sweden External Links: Link Cited by: §1.1, Table 2.
  • M. Sekoyan, N. R. Koluguri, N. Tadevosyan, P. Zelasko, T. Bartley, N. Karpov, J. Balam, and B. Ginsburg (2025) Canary-1b-v2 & parakeet-tdt-0.6b-v3: efficient and high-performance models for multilingual asr and ast. External Links: 2509.14128, Link Cited by: §6.1.
  • P. E. Solberg and P. Ortiz (2022) The Norwegian parliamentary speech corpus. In Proceedings of the Thirteenth Language Resources and Evaluation Conference, Marseille, France, pp. 1003–1008. Cited by: §1.1, Table 2.
@@
  • V. Srivastav, S. Majumdar, N. Koluguri, A. Moumen, S. Gandhi, et al. (2023) Open automatic speech recognition leaderboard. Hugging Face. External Links: Link Cited by: §1.
  • V. Stankov, M. Kopp, and O. Bojar (2025) ParCzech4Speech: a new speech corpus derived from Czech parliamentary data. In Proceedings of the 28th International Conference on Text, Speech, and Dialogue, pp. to appear. Cited by: §1.1, Table 2.
@@
  • SYSTRAN (2025) SYSTRAN/faster-whisper Note: GitHub repository External Links: Link Cited by: §9.
  • V. Timmel, C. Paonessa, R. Kakooee, M. Vogel, and D. Perruchoud (2024) Fine-tuning Whisper on low-resource languages for real-world applications. arXiv preprint arXiv:2412.15726. Cited by: §1.
  • A. Virkkunen, A. Rouhe, N. Phan, and M. Kurimo (2023) Finnish parliament ASR corpus: analysis, benchmarks and statistics. Language Resources and Evaluation 57 (4), pp. 1645–1670. Cited by: §1.1, Table 2.
@@
  • C. Wang, M. Riviere, A. Lee, A. Wu, C. Talnikar, D. Haziza, M. Williamson, J. Pino, and E. Dupoux (2021) VoxPopuli: a large-scale multilingual speech corpus for representation learning, semi-supervised learning and interpretation. arXiv preprint arXiv:2101.00390. Cited by: §1.2, §3.
@@
  • T. Wolf, L. Debut, V. Sanh, J. Chaumond, C. Delangue, A. Moi, P. Cistac, T. Rault, R. Louf, M. Funtowicz, J. Davison, S. Shleifer, P. von Platen, C. Ma, Y. Jernite, J. Plu, C. Xu, T. L. Scao, S. Gugger, M. Drame, Q. Lhoest, and A. M. Rush (2020) Transformers: state-of-the-art natural language processing. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Online, pp. 38–45. External Links: Link Cited by: §5.3.1.
@@
  • Y. Zhao, A. Gu, R. Varma, L. Luo, C. Huang, M. Xu, L. Wright, H. Shojanazeri, M. Ott, S. Shleifer, et al. (2023) Pytorch fsdp: experiences on scaling fully sharded data parallel. arXiv preprint arXiv:2304.11277. Cited by: §5.3.1.